# On the mixed UDA states and additivity

### v1

@@ I Introduction
Understanding the information of a global state with its local parts elucidates profound differences between quantum mechanics and its classical counterpart. Quantum marginal problem deals with the existence of a valid global state, whose marginals (i.e., reduced density matrices) match the input set of states coleman1963structure; klyachko2004quantum. If such a quantum marginal problem does admit a solution, one may focus on an interrelated problem. Given access to local information provided by the marginals, is the global state determined uniquely by the marginals? If yes, this global state is uniquely determined among all (UDA) states, which naturally follows the need of quantum system certification, including quantum state tomography (QST) bisio2009optimal; huang2025process, self-testing ivan2023quantum, and other device-independent tasks qiu2024measurement. Generally, the number of observables required for certification increases exponentially with the number of systems qin2024state, whereas the UDA property allows the required observables exhibits polynomial growth. This lowers the complexity of reconstructing the state of systems substantially cramer2010efficient; experiment2024hu. Besides, the UDA property brings tools for dissipative quantum control karuvade2019uniquely, and topological stabilizer codes in multipartite quantum error correction liao2022topo. Moreover, the characterization of a set of marginals that uniquely determines the global state sheds light on the structure of multipartite entanglement across different subsystems den2002almost; walck2008only.
@@
The UDA problem arises from the scope of pure states. Given a pure state $|\psi\rangle$, if there are no other states sharing the same marginals with $|\psi\rangle$, it is said to be uniquely determined. This is further clarified into two categories: uniquely determined among pure (UDP) states and uniquely determined among all (UDA) states. By the two categories, pure states can be neither UDP nor UDA, UDP but not UDA, or UDA states qt2017tao. The characterization of pure, uniquely determined states has been extensively studied. Almost all three-qubit pure states, general symmetric Dicke and $W$-type states have been proved to be UDA by 2-marginals den2002almost; rana2011optimal; yu2013multi; zhang2024EvenParticle. The only $n$-particle pure states which can not be UDP by their (n1)(n-1)-marginals are those GHZ-type states, and the result is further improved to the case of UDA walck2008only. In addition to the pure states, mixed states can also be uniquely determined among all states, for example, a tripartite UDA state with equal-dimensional subsystems has been constructed role2014chen. Mixed states exist commonly in laboratories than pure ones since real quantum systems inevitably interact with their environment, and such decoherence and noise destroy purity liu2023observation; lee2025topo. From this perspective, the characterization of mixed UDA states has both fundamental and practical implications. Compared with pure states, UDA mixed states have been less explored. The certification of mixed UDA states is not easy, even when the state is given numerically. It is shown that almost every tripartite density operator with low rank is UDA by two specific bipartite marginals, and thus almost all four-qubit states of rank two are UDA by tripartite marginals uinque2013chen. For the convenience of simplification, the scope of some existing results is focused on ”almost all states”, by excluding a measure-zero set of states jones2005parts; wyder2017almost; parashar2009nqubit. However, for a state under consideration, it is not straightforward to verify whether it is in the measure-zero set. Therefore, it is of practical interest to give a complete characterization of mixed UDA states by giving the necessary and sufficient conditions. This is the first motivation of this work.
@@
For a set of marginals on subsystems of fixed size, the global state may transition from a UDA state to a non-UDA state as the number of systems increases. For example, generic states of $n$-qudits are UDA by certain sets of marginals of less than about two-thirds linden2002parts, or more than half of the parties jones2005parts, whereas the reduced states of fewer than half of the parties are not sufficient. This implies that given two mixed UDA states $\alpha$ and $\beta$, their additive form $\alpha\otimes\beta$ is not necessarily UDA. Admitting that one of $\alpha$ or $\beta$ is pure, $\alpha\otimes\beta$ is shown to be UDA by three composite types of tensor product add2023shen. Generally, it is worth considering the additivity of two mixed UDA states by focusing on the following questions: can the UDA additivity hold without additional conditions? If the answer is negative, which kind of additional conditions of $\alpha$ and $\beta$ make the UDA additivity hold? The purity assumption in add2023shen avoids analyzing the essence of UDA states, whereas this is unavoidable when both states are mixed. By fully characterizing mixed UDA states, the additivity of these UDA states can be further studied. This is the second motivation of this work.
@@
In this work, we study the complete characterization of multipartite mixed UDA states, and show the necessary and sufficient condition under which the additivity of UDA states holds. As the beginning, we introduce some essential properties for arbitrary UDA states in Lemmas 4-7. In detail, we show that the global state is not UDA if its marginals or faces of the range are not UDA, and that two states with the same range, or up to permutations, are either UDA or not at the same time. We also show the range inclusion between a specific state and its compatible states. By these facts, we show the complete characterization of bipartite 1-UDA states in Proposition 8. The necessary and sufficient conditions by which some three- to $n$-partite states are UDA are given in Theorems 14, 20, 22 and Proposition 23. We clarify that the study of $k$-UDA, particularly for small $k$, is motivated by practical considerations. Since the measurement cost, the error of gates and the impact of noise grow with $k$, minimizing resource consumption in local measurements makes $2$-UDA a favorable scenario. Inspired by the proof of preceding facts, a systematic method is derived for the determination of $k$-UDA states in FIG. 1. The UDA additivity of bipartite, three-qubit 2-UDA and some $n$-partite states is further analyzed. For the bipartite UDA states $\alpha$ and $\beta$, $\alpha\otimes\beta$ is UDA if and only if one of $\alpha$ and $\beta$ is a pure product state, shown in Theorem 26. The additivity of three-qubit UDA states is given in Theorem 27, by showing the necessary and sufficient condition. The applications of mixed UDA states and their characterization are shown from the perspectives of tomography, entanglement certification and other tasks.
@@ II Preliminaries
The existence of mixed $k$-UDA states is established in previous work. It is known that some mixed states with low rank can be UDA. For example, Theorem 4 in uinque2013chen shows that almost every tripartite density operator $\rho$ acting on the Hilbert space ${\cal H}_{d_{1}}\otimes{\cal H}_{d_{2}}\otimes{\cal H}_{d_{3}}$ with rank no more than $\lfloor\frac{d_{1}}{d_{3}}\rfloor$ can be uniquely determined among all states by its 2-marginals of partial pairs $\{1,2\}$ and $\{1,3\}$. It is also shown in uinque2013chen that giving a complete characterization of these undetermined tripartite states is not easy, as the methods proposed before do not directly provide it. This inspires us to give a complete characterization of UDA states before analyzing their additivity. That is, we establish a criterion to determine whether a mixed state is UDA directly, which is exactly what we will show in Sec. III.
A state with fewer parties may transform from a $k$-UDA state into a non-$k$-UDA one, with the increase of the number of parties with a fixed $k\in\mathbb{N}$. A construction of $k$-UDA states in a system with more parties is considered in add2023shen. In detail, three composite types of tensor product for two $k$-UDA states preserve $k$-UDA if one of the two $k$-UDA states is pure. This relies on the following fact, which we shall use later.
@@
Obviously, Lemma 2 does not generally hold if the $m$-partite marginal $\rho_{A_{1}\cdots A_{m}}$ is a mixed state. For example, the state $\rho_{ABC}=p|000\rangle\!\langle 000|+(1-p)|W\rangle\!\langle W|$ with $|W\rangle=\frac{1}{\sqrt{3}}(|000\rangle+|001\rangle+|100\rangle)$ satisfies that $\rho_{AB}$ and $\rho_{C}$ are both mixed, and $\rho_{ABC}\neq\rho_{AB}\otimes\rho_{C}$. Hence, the construction in add2023shen fails if both initial states are mixed. It is interesting to establish additional conditions on the initial states, by which the additivity of mixed UDA states still holds.
We introduce some basic properties for a $k$-UDA state add2023shen as follows. Let $\rho$ be a $n$-partite $k$-UDA state. Then (i) any state that is LU equivalent to $\rho$ is also a $k$-UDA state; (ii) $\rho$ is also $(k+1)$-UDA state for $k+1\leq n$, and the converse does not generally hold. These facts allow us to simplify the form of density operators in the analysis of their UDA properties, which shall be widely used in Sec. III. In addition to the equation of $k$-marginals, the positive semi-definiteness is a natural restriction of a compatible state. To prove that some states are not $k$-UDA, we construct another state compatible with them, and the semi-definiteness of these states is guaranteed by the following fact.
@@ III.1 Preliminary lemmas
We show some essential properties satisfied by arbitrary $k$-UDA states from the perspective of their ”components” and range. We also show that the UDA property is invariant to the permutation operations. First, Lemma 4 and 5 show that if the marginals and the element of the range of a state is not $k$-UDA, then the whole state is also not $k$-UDA, respectively. Then we obtain that two states that have the same range are the same time $k$-UDA or not. Lemma 6 shows that the states compatible with each other satisfy range inclusion. Besides, Lemma 7 shows that the UDA property is invariant up to system permutation. The proof of Lemmas 4-7 is given in Appendix A.
@@
The preceding facts contribute to simplifying the analysis of UDA properties by applying permutation operations. In this paper, we shall interchange the order of any two subsystems of a product state without explicit clarification.
@@ III.3 On tripartite mixed UDA states
Let $\rho=\rho_{A_{1}}\otimes\rho_{A_{2}A_{3}}$ be a tripartite state supported on $\mathbb{C}^{d_{1}}\otimes\mathbb{C}^{d_{2}}\otimes\mathbb{C}^{d_{3}}$. Suppose $\mathop{\rm rank}(\rho_{A_{1}})=d_{1}$. Then
(i) If $\mathop{\rm rank}(\rho_{A_{2}A_{3}})=d_{2}d_{3}-1$ and $\rho_{A_{2}A_{3}}$ is not LU equivalent to the state with a zero diagonal element, then $\rho$ is not 2-UDA.
(ii) If $\mathop{\rm rank}(\rho_{A_{2}A_{3}})=d_{2}d_{3}$, then $\rho$ is not 2-UDA.
@@
By summarizing the facts in Propositions 12 and 13, we obtain the complete characterization of a 2-UDA product state $\rho_{A_{1}}\otimes\rho_{A_{2}A_{3}}$ as follows.
@@
Based on Theorem 14, the whole three-qubit mixed states that satisfy one of the three conditions can be determined by their 2-marginals. This implies that it is practically redundant to perform tomography on the system $A_{1}A_{2}A_{3}$.
@@ III.4 On four-qubit mixed UDA states
By now, we have completed the analysis of the 2-UDA property of any four-qubit state $\rho=\rho_{A_{1}A_{2}}\otimes\rho_{A_{3}A_{4}}$. We conclude the results of Propositions 17, 18, and 19, and obtain the necessary and sufficient condition by which $\rho$ is 2-UDA.
@@ III.5 On $n$-qubit mixed UDA states for $n\geq 5$
The proof of Lemma 21 is given in Appendix D. Using Theorems 14 and 20, Proposition 23 and Lemma 21, we have the complete characterization of a five-qubit product state.
@@
The proof of Proposition 23 is given in Appendix D. We shall emphasize that a $k$-UDA state is also $(k+1)$-UDA. This implies that $\rho=\otimes_{i=1}^{n-2}\rho_{A_{i}}\otimes\rho_{A_{n-1}A_{n}}$ is $k$-UDA if the conditions in Proposition 23 (ii) are satisfied.
@@
(i) Every full-rank $n$-partite mixed state supported on $\mathbb{C}^{d_{1}}\otimes\mathbb{C}^{d_{2}}\otimes...\otimes\mathbb{C}^{d_{n}}$ is not $k$-UDA, for $1\leq k\leq n-1$.
(ii) Every $n$-qubit mixed state of rank $2^{n}-1$ is not 2-UDA.
@@
(i) If $\rho$ is full-rank, then the range ${\cal R}(\rho)$ includes the quasi-GHZ state $a|0\rangle^{\otimes n}+b|1\rangle^{\otimes n}$, which is not $k$-UDA. Thus $\rho$ is not 2-UDA.
(ii) Let $\rho$ be an $n$-qubit mixed state. Suppose $\mathop{\rm rank}\rho=2^{n}-1$. We first prove that $\rho$ is not 2-UDA for $n=3$. We claim that the range of $\rho$ has a pure state $|\gamma\rangle$ LU equivalent to the quasi-GHZ state $a|000\rangle+b|111\rangle$ with $ab\neq 0$. Up to LU equivalence, the kernel of a seven-dimensional space $V\subseteq\mathbb{C}^{8}$ satisfies that $V^{\perp}=\text{span}\{a_{0}|000\rangle+a_{1}|001\rangle+a_{2}|010\rangle+e^{i\theta}a_{3}|011\rangle+a_{4}|111\rangle\}$ acin2000generalize. The state $|\beta\rangle=a_{2}^{*}|001\rangle+\frac{a_{1}^{*}a_{2}}{x^{*}}|110\rangle-a_{1}^{*}|010\rangle+x|101\rangle$ is LU equivalent to the quasi-GHZ state and $|\beta\rangle\perp V^{\perp}$. Therefore, the non-2-UDA state $|\beta\rangle\in V$.
It remains to prove the claim for $n$-qubit states $\rho$ of rank $2^{n}-1$ and $n>3$. It suffices to show that the range of $\rho$ has a pure state LU to the state $|\gamma\rangle|0\rangle^{\otimes(n-3)}$. Let $\ker\rho$ be spanned by the pure state $|\psi\rangle=\sum_{(j_{1},...,j_{n-3})\neq(0,...,0)}|\alpha_{j_{1}...j_{n-3}}\rangle|j_{1}...j_{n-3}\rangle+|\beta\rangle|0\rangle^{\otimes n-3}$. Using the proof for $n=3$, we can assume that $|\beta\rangle$ is orthogonal to $|\gamma\rangle$. Hence $|\psi\rangle$ is orthogonal to $|\gamma\rangle|0\rangle^{\otimes(n-3)}$. We have proven the claim.   $\sqcap\sqcup$
With respect to any unitarily invariant induced measure $\mu_{N,K}$ with $K\geq N$ (in particular the Hilbert–Schmidt or Bures measure), the set of rank-deficient density matrices has measure zero. Indeed, the joint eigenvalue density Zyczkowski2001induced is absolutely continuous on the eigenvalue simplex and vanishes on its boundary. This implies the following fact.
@@
It is shown that pure states are typically UDA, which allows local quantum state tomography qin2024state. Corollary 25 shows that the UDA property is unavoidable in the certification of mixed states by such kind of tomography.

### v2

@@ I Introduction
Understanding the information of a global state with its local parts elucidates profound differences between quantum mechanics and its classical counterpart. Quantum marginal problem deals with the existence of a valid global state, whose marginals (i.e., reduced density matrices) match the input set of states Coleman (1963); Klyachko . If such a quantum marginal problem does admit a solution, one may focus on an interrelated problem. Given access to local information provided by the marginals, is the global state determined uniquely by the marginals? If yes, this global state is uniquely determined among all (UDA) states, which naturally follows the need of quantum system certification, including quantum state tomography (QST) Bisio et al. (2009); Huang et al. (2025b), self-testing Šupić et al. (2023), and other device-independent tasks Qiu and Chen (2024). Generally, the number of observables required for certification increases exponentially with the number of systems Qin et al. (2024), whereas the UDA property allows the required observables exhibits polynomial growth. This lowers the complexity of reconstructing the state of systems substantially Cramer et al. (2010); Hu et al. (2024). Besides, the UDA property brings tools for dissipative quantum control Karuvade et al. (2019), and topological stabilizer codes in multipartite quantum error correction Liao et al. (2022). Moreover, the characterization of a set of marginals that uniquely determines the global state sheds light on the structure of multipartite entanglement across different subsystems Linden et al. (2002); Walck and Lyons (2008).
@@
The UDA problem arises from the scope of pure states. Given a pure state $|\psi\rangle$, if there are no other states sharing the same marginals with $|\psi\rangle$, it is said to be uniquely determined. This is further clarified into two categories: uniquely determined among pure (UDP) states and uniquely determined among all (UDA) states. By the two categories, pure states can be neither UDP nor UDA, UDP but not UDA, or UDA states Xin et al. (2017). The characterization of pure, uniquely determined states has been extensively studied. Almost all three-qubit pure states, general symmetric Dicke and $W$-type states have been proved to be UDA by 2-marginals Linden et al. (2002); Rana and Parashar (2011); Yu (2013); Zhang et al. (2024). The only $n$-particle pure states which can not be UDP by their (n1)(n-1)-marginals are those GHZ-type states, and the result is further improved to the case of UDA Walck and Lyons (2008). In addition to the pure states, mixed states can also be uniquely determined among all states, for example, a tripartite UDA state with equal-dimensional subsystems has been constructed Chen et al. (2014). Mixed states exist commonly in laboratories than pure ones since real quantum systems inevitably interact with their environment, and such decoherence and noise destroy purity Liu et al. (2023); Lee and Moon (2025). From this perspective, the characterization of mixed UDA states has both fundamental and practical implications. Compared with pure states, UDA mixed states have been less explored. The certification of mixed UDA states is not easy, even when the state is given numerically. It is shown that almost every tripartite density operator with low rank is UDA by two specific bipartite marginals, and thus almost all four-qubit states of rank two are UDA by tripartite marginals Chen et al. (2013). For the convenience of simplification, the scope of some existing results is focused on ”almost all states”, by excluding a measure-zero set of states Jones and Linden (2005); Wyderka et al. (2017); Parashar and Rana (2009). However, for a state under consideration, it is not straightforward to verify whether it is in the measure-zero set. Therefore, it is of practical interest to give a complete characterization of mixed UDA states by giving the necessary and sufficient conditions. This is the first motivation of this work.
@@
For a set of marginals on subsystems of fixed size, the global state may transition from a UDA state to a non-UDA state as the number of systems increases. For example, generic states of $n$-qudits are UDA by certain sets of marginals of less than about two-thirds Linden and Wootters (2002), or more than half of the parties Jones and Linden (2005), whereas the reduced states of fewer than half of the parties are not sufficient. This implies that given two mixed UDA states $\alpha$ and $\beta$, their additive form $\alpha\otimes\beta$ is not necessarily UDA. Admitting that one of $\alpha$ or $\beta$ is pure, $\alpha\otimes\beta$ is shown to be UDA by three composite types of tensor product Shen and Chen (2023). Generally, it is worth considering the additivity of two mixed UDA states by focusing on the following questions: can the UDA additivity hold without additional conditions? If the answer is negative, which kind of additional conditions of $\alpha$ and $\beta$ make the UDA additivity hold? The purity assumption in Shen and Chen (2023) avoids analyzing the essence of UDA states, whereas this is unavoidable when both states are mixed. By fully characterizing mixed UDA states, the additivity of these UDA states can be further studied. This is the second motivation of this work.
@@
In this work, we study the complete characterization of multipartite mixed UDA states, and show the necessary and sufficient condition under which the additivity of UDA states holds. As the beginning, we introduce some essential properties for arbitrary UDA states in Lemmas 4-8. In detail, we show that the global state is not UDA if its marginals or faces of the range are not UDA, and that two states with the same range, or up to permutations and reversible channels, are either UDA or not at the same time. We also show the range inclusion between a specific state and its compatible states, and the real and imaginary parts of a state. By these facts, we show the complete characterization of bipartite 1-UDA states in Proposition 9. The necessary and sufficient conditions by which some three- to $n$-partite states are UDA are given in Theorems 15, 22, 24 and Proposition 25. We clarify that the study of $k$-UDA, particularly for small $k$, is motivated by practical considerations. Since the measurement cost, the error of gates and the impact of noise grow with $k$, minimizing resource consumption in local measurements makes $2$-UDA a favorable scenario. Inspired by the proof of preceding facts, a systematic method is derived for the determination of $k$-UDA states in FIG. 1. The UDA additivity of bipartite, three-qubit 2-UDA and some $n$-partite states is further analyzed. For the bipartite UDA states $\alpha$ and $\beta$, $\alpha\otimes\beta$ is UDA if and only if one of $\alpha$ and $\beta$ is a pure product state, shown in Theorem 28. The additivity of three-qubit UDA states is given in Theorem 29, by showing the necessary and sufficient condition. The applications of mixed UDA states and their characterization are shown from the perspectives of tomography, entanglement certification and other tasks.
@@ II Preliminaries
The existence of mixed $k$-UDA states is established in previous work. It is known that some mixed states with low rank can be UDA. For example, Theorem 4 in Chen et al. (2013) shows that almost every tripartite density operator $\rho$ acting on the Hilbert space ${\cal H}_{d_{1}}\otimes{\cal H}_{d_{2}}\otimes{\cal H}_{d_{3}}$ with rank no more than $\lfloor\frac{d_{1}}{d_{3}}\rfloor$ can be uniquely determined among all states by its 2-marginals of partial pairs $\{1,2\}$ and $\{1,3\}$. It is also shown in Chen et al. (2013) that giving a complete characterization of these undetermined tripartite states is not easy, as the methods proposed before do not directly provide it. This inspires us to give a complete characterization of UDA states before analyzing their additivity. That is, we establish a criterion to determine whether a mixed state is UDA directly, which is exactly what we will show in Sec. III.
A state with fewer parties may transform from a $k$-UDA state into a non-$k$-UDA one, with the increase of the number of parties with a fixed $k\in\mathbb{N}$. A construction of $k$-UDA states in a system with more parties is considered in Shen and Chen (2023). In detail, three composite types of tensor product for two $k$-UDA states preserve $k$-UDA if one of the two $k$-UDA states is pure. This relies on the following fact, which we shall use later.
@@
Obviously, Lemma 2 does not generally hold if the $m$-partite marginal $\rho_{A_{1}\cdots A_{m}}$ is a mixed state. For example, the state $\rho_{ABC}=p|000\rangle\!\langle 000|+(1-p)|W\rangle\!\langle W|$ with $|W\rangle=\frac{1}{\sqrt{3}}(|000\rangle+|001\rangle+|100\rangle)$ satisfies that $\rho_{AB}$ and $\rho_{C}$ are both mixed, and $\rho_{ABC}\neq\rho_{AB}\otimes\rho_{C}$. Hence, the construction in Shen and Chen (2023) fails if both initial states are mixed. It is interesting to establish additional conditions on the initial states, by which the additivity of mixed UDA states still holds.
We introduce some basic properties for a $k$-UDA state Shen and Chen (2023) as follows. Let $\rho$ be an $n$-partite $k$-UDA state. Then (i) any state that is LU equivalent to $\rho$ is also a $k$-UDA state; (ii) $\rho$ is also $(k+1)$-UDA state for $k+1\leq n$, and the converse does not generally hold. These facts allow us to simplify the form of density operators in the analysis of their UDA properties, which shall be widely used in Sec. III. In addition to the equation of $k$-marginals, the positive semi-definiteness is a natural restriction of a compatible state. To prove that some states are not $k$-UDA, we construct another state compatible with them, and the semi-definiteness of these states is guaranteed by the following fact.
@@ III.1 Preliminary lemmas
We show some essential properties satisfied by arbitrary $k$-UDA states from the perspective of their ”components” and range. We also show that the UDA property is invariant to the permutation operations. First, Lemma 4 and 5 show that if the marginals and the element of the range of a state is not $k$-UDA, then the whole state is also not $k$-UDA, respectively. Then we obtain that two states that have the same range are the same time $k$-UDA or not. Lemma 6 shows that the states compatible with each other satisfy range inclusion. Besides, Lemma 7 shows that the UDA property is invariant up to system permutation. The proof of Lemmas 4-8 is given in Appendix A.
@@
The proof is given in Appendix A. The preceding facts contribute to simplifying the analysis of UDA properties by applying permutation operations. In this paper, we shall interchange the order of any two subsystems of a product state without explicit clarification.
From a general perspective, we show the properties for $k$-UDA states and any two compatible states.
Lemma 8
(i) If two multipartite states $\rho$ and $\sigma$ are $k$-compatible, then their real and imaginary parts are also $k$-compatible.
(ii) Let $\rho=\text{Re}(\rho)+i\text{Im}(\rho)$. Then ${\cal R}(\text{Im}(\rho))\subseteq{\cal R}(\text{Re}(\rho))$.
(iii) Suppose $\rho$ is not a real state, and all the $k$-marginals of $\rho$ are real. Then $\rho$ is not $k$-UDA.
(iv) Suppose the $n$-partite state $\rho\geq 0$ and $H$ is an Hermitian matrix of the same size, and ${\cal R}(\rho)\supseteq{\cal R}(H)$. Then ${\cal R}(\rho_{\cal S})\supseteq{\cal R}(H_{\cal S})$ for ${\cal S}\subseteq[A_{1}A_{2}...A_{n}]$. The converse does not generally hold.
(v) Suppose $\alpha$ and $\beta$ are $k$-compatible $n$-partite states on systems $A_{1}A_{2}...A_{n}]$, and $\Lambda$ a channel on the first system $A_{1}$, with the remaining systems denoted as $B:=A_{2}...A_{n}$. Then (a) $(\Lambda_{A_{1}}\otimes I_{B})\alpha$ and $(\Lambda_{A_{1}}\otimes I_{B})\beta$ are also $k$-compatible; (b) if $\Lambda$ is a reversible channel, then $\alpha$ is $k$-UDA iff $(\Lambda_{A_{1}}\otimes I_{B})\alpha$ is $k$-UDA.
The proof is given in Appendix A. These facts provide the tools to develop the results for UDA states from the perspective of the real and image parts, and their range. Lemma 8 (v) generalizes the fact that UDA property is invariant up to LU equivalence. The conditions of reversible channel and Petz recovery maps JENCOVA (2012) may contribute to the analyze of arbitrary multipartite UDA states.
@@ III.3 On tripartite mixed UDA states
Let $\rho=\rho_{A_{1}}\otimes\rho_{A_{2}A_{3}}$ be a tripartite state supported on $\mathbb{C}^{d_{1}}\otimes\mathbb{C}^{2}\otimes\mathbb{C}^{2}$. Suppose $\mathop{\rm rank}(\rho_{A_{1}})=d_{1}$. Then
(i) If $\mathop{\rm rank}(\rho_{A_{2}A_{3}})=3$ and $\rho_{A_{2}A_{3}}$ is not LU equivalent to the state with a zero diagonal element, then $\rho$ is not 2-UDA.
(ii) If $\mathop{\rm rank}(\rho_{A_{2}A_{3}})=4$, then $\rho$ is not 2-UDA.
@@
By summarizing the facts in Propositions 13 and 14, we obtain the complete characterization of a 2-UDA product state $\rho_{A_{1}}\otimes\rho_{A_{2}A_{3}}$ as follows.
@@
The following fact shows that a general three-qubit state with high rank is not 2-UDA, which is a special case of Proposition 26.
Lemma 16
A three-qubit state $\rho_{A_{1}A_{2}A_{3}}$ with $\mathop{\rm rank}(\rho)\geq 7$ is not 2-UDA. Further, if $\mathop{\rm rank}(\rho)=6$ and $\mathop{\rm rank}(\rho_{A_{2}A_{3}})=3$, $\rho$ is not 2-UDA.
The proof is given in Appendix B. The complete characterization of a general mixed state remains to be further explored.
@@ III.4 On four-qubit mixed UDA states
By now, we have completed the analysis of the 2-UDA property of any four-qubit state $\rho=\rho_{A_{1}A_{2}}\otimes\rho_{A_{3}A_{4}}$. We conclude the results of Propositions 19, 20, and 21, and obtain the necessary and sufficient condition by which $\rho$ is 2-UDA.
@@ III.5 On $n$-qubit mixed UDA states for $n\geq 5$
The proof of Lemma 23 is given in Appendix D. Using Theorems 15 and 22, Proposition 25 and Lemma 23, we have the complete characterization of a five-qubit product state.
@@
The proof of Proposition 25 is given in Appendix D. We shall emphasize that a $k$-UDA state is also $(k+1)$-UDA. This implies that $\rho=\otimes_{i=1}^{n-2}\rho_{A_{i}}\otimes\rho_{A_{n-1}A_{n}}$ is $k$-UDA if the conditions in Proposition 25 (ii) are satisfied. Further, we introduce some $n$-qubit 2-UDA states. An $n$-qubit state $\rho$ is a symmetric state if $\rho=\text{Swap}_{(i,j)}\rho$ for all $1\leq i,j\leq n$. It is shown that if all 2-marginals of an $n$-qubit state $\rho$ are symmetric, then $\rho$ is symmetric Shi et al. (2025). Actually, we only need n1n-1 2-marginals corresponding to a connected graph, say $\rho_{12},\rho_{23},...,\rho_{(n-1)n}$. Hence, the multiqubit Dicke state is 2-UDA.
@@
(i) Every $n$-partite mixed state $\rho$ supported on $\mathbb{C}^{d_{1}}\otimes\mathbb{C}^{d_{2}}\otimes...\otimes\mathbb{C}^{d_{n}}$ with $\mathop{\rm rank}(\rho)\geq d_{1}d_{2}...d_{n}-1$ is not $k$-UDA, for $1\leq k\leq n-1$. Besides, the full-rank $\rho$ can be compatible with another full rank state.
(ii) The $n$-qubit state $\rho_{A_{1}A_{2}...A_{n}}$ with $\mathop{\rm rank}(\rho)=d_{1}(d_{2}d_{3}...d_{n}-1)$ and $\mathop{\rm rank}(\rho_{A_{2}...A_{n}})=d_{2}d_{3}...d_{n}-1$, then the state is not $k$-UDA, for $1\leq k\leq n-2$.
@@
(i) For the full-rank $\rho$, we can choose a nonzero Hermitian operator $\chi$ satisfying $\mathop{\rm Tr}_{A_{i_{1}}...A_{i_{k}}}(\chi)=0$, $\forall i_{j}\in\{1,2,...,n\}$. We further choose $\sigma=\rho+\epsilon\chi$, where $\epsilon$ is small enough to guarantee that $\sigma>0$. Thus $\rho$ is not $k$-UDA and compatible with another full-rank state $\sigma$.
Suppose $\mathop{\rm rank}(\rho)=d_{1}d_{2}...d_{n}-1$. Firstly, we consider the $n$-qubit state $\rho$ for $\mathop{\rm rank}\rho=2^{n}-1$. We prove that $\rho$ is not 2-UDA for $n=3$. We claim that the range of $\rho$ has a pure state $|\gamma\rangle$ LU equivalent to the quasi-GHZ state $a|000\rangle+b|111\rangle$ with $ab\neq 0$. Up to LU equivalence, the kernel of a seven-dimensional space $V\subseteq\mathbb{C}^{8}$ satisfies that $V^{\perp}=\text{span}\{a_{0}|000\rangle+a_{1}|001\rangle+a_{2}|010\rangle+e^{i\theta}a_{3}|011\rangle+a_{4}|111\rangle\}$ Acin et al. (2000). The state $|\beta\rangle=a_{2}^{*}|001\rangle+\frac{a_{1}^{*}a_{2}}{x^{*}}|110\rangle-a_{1}^{*}|010\rangle+x|101\rangle$ is LU equivalent to the quasi-GHZ state and $|\beta\rangle\perp V^{\perp}$. Therefore, the non-2-UDA state $|\beta\rangle\in V$. It remains to prove the claim for $n$-qubit states $\rho$ of rank $2^{n}-1$ and $n>3$. It suffices to show that the range of $\rho$ has a pure state LU to the state $|\gamma\rangle|0\rangle^{\otimes(n-3)}$. Let $\ker\rho$ be spanned by the pure state $|\psi\rangle=\sum_{(j_{1},...,j_{n-3})\neq(0,...,0)}|\alpha_{j_{1}...j_{n-3}}\rangle|j_{1}...j_{n-3}\rangle+|\beta\rangle|0\rangle^{\otimes n-3}$. Using the proof for $n=3$, we can assume that $|\beta\rangle$ is orthogonal to $|\gamma\rangle$. Hence $|\psi\rangle$ is orthogonal to $|\gamma\rangle|0\rangle^{\otimes(n-3)}$.
Next we consider the case for $\mathop{\rm rank}(\rho)=d_{1}d_{2}...d_{n}-1$. Let $|\psi\rangle$ be a state supported on $\otimes_{k=1}^{n}\mathbb{C}^{d_{k}}$. Then $|\psi\rangle=(P+Q)|\psi\rangle$, where $Q$ is the projector onto the $n$-qubit space $H_{Q}$, and $P+Q=I_{\prod_{j}d_{j}}$. Let $S=\text{span}\{Q|\psi\rangle\}^{\perp}$ regarding to $H_{Q}$. We have $\mathop{\rm Dim}S=2^{n}-1$, and $\forall|x\rangle\in S$, $\langle x|\psi\rangle=0$, i.e. $S=\text{span}\{|\psi\rangle\}^{\perp}$. By taking $|\psi\rangle$ as the vector orthogonal to ${\cal R}(\rho)$, we have $S\subseteq{\cal R}(\rho)$. By the preceding analysis, $S$ contains a pure state LU equivalent to the quasi-GHZ state. Hence $\rho$ is not $k$-UDA by Lemma 5 (i). We have proven the claim.
(ii) Similar to the proof of Lemma 16, the range of $\rho_{A_{1}...A_{n}}$ is the same as that of $\rho_{A_{1}}\otimes\rho_{A_{2}A_{3}...A_{n}}$. By the result in (i), $\rho_{A_{2}A_{3}...A_{n}}$ is not $k$-UDA for $1\leq k\leq n-2$. Using Lemma 4 and 5, $\rho_{A_{1}...A_{n}}$ is not $k$-UDA.   $\sqcap\sqcup$
With respect to any unitarily invariant induced measure $\mu_{N,K}$ with $K\geq N$ (in particular the Hilbert–Schmidt or Bures measure), the set of rank-deficient density matrices has measure zero. Indeed, the joint eigenvalue density Życzkowski and Sommers (2001) is absolutely continuous on the eigenvalue simplex and vanishes on its boundary. This implies the following fact.
@@
It is shown that pure states are typically UDA, which allows local quantum state tomography Qin et al. (2024). Corollary 27 shows that the UDA property is unavoidable in the certification of mixed states by such kind of tomography.